OpenStudy (anonymous):

what is the slope of the line that contains the points (-3,2) and (6,-2) ?

OpenStudy (the_fizicx99):

Same thing: M = -2-2/6-(-3)

OpenStudy (anonymous):

?

OpenStudy (the_fizicx99):

Subtract that

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(-3\quad ,&2)\quad &(6\quad ,&-2) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

I got a decimal ... we i subtracted all of that.

OpenStudy (anonymous):

when

OpenStudy (jdoe0001):

yes.. you can use a decimal, usually you'd leave it as fraction though

OpenStudy (the_fizicx99):

-2-2 = -4 6+3 = 9 -4/9

OpenStudy (the_fizicx99):

As a fraction because you use it as rise/run

OpenStudy (anonymous):

thanks fizi. ill give u a medal :)

OpenStudy (the_fizicx99):

oki

OpenStudy (anonymous):

and i became ur fan